Of the $50,000 that Natasha pocketed on her last real estate deal, $20,000 went to charity. She invested part of the remainder in Dreyfus New Leaders Fund with an annual yield of 16% and the rest in Templeton Growth Fund with an annual yield of 25%. If she made $6060 on these investments in 1 year, then how much did she invest in each fund?
Answer by mananth(16949)

$ 50,000 was the amount received.
20000 went to charity
So she invested $30,000
Let her invest x in New Leader - 16%
she will invest 30,000-x in Templeton 25%
..
Total she earned =
$9090
.
0.16x+0.25(30,000-x)=6060
0.16x+7500-0.25x=6060
-0.09x=-1440
x= -1440 / -0.09
x= $ 16,000 with Dreyfus New leaders
$14,000 with Templeton